OpenStudy (anonymous):

Prove the statement using the epsilon and delta definition of a limit. lim as x -> 3 (x^2 +x -4) = 8

OpenStudy (anonymous):

we can do this. ready?

OpenStudy (anonymous):

first of all like in all these problems we work backwards

OpenStudy (anonymous):

okay =)

OpenStudy (anonymous):

we want a delta, but first we see how to force \[x^2+x-4\] to be close to 8, but which we mean \[|x^2+x-4-8|<\epsilon\]

OpenStudy (anonymous):

so we are really looking at \[|x^2+x-12|<\epsilon\] and we want a delta that will make this true if \[|x-3|<\delta\]

OpenStudy (anonymous):

first we factor and see \[|x^2+x-12|<\epsilon\] is the same thing as \[|(x-3)(x+4)|<\epsilon\] and now this is not too bad because we have control over the \[(x-3)\] factor. unfortunately we also have \[(x+4)\] to contend with.

OpenStudy (anonymous):

so the gimmick here is to say that lets make sure that \[|x-3|<1\] say. this will give a bound on \[|x+4|\] don't forget we get to control \[|x-3|\] so we can make it as small as we like and we can certainly make it smaller than 1

OpenStudy (anonymous):

now if \[|x-3|<1\] that forces \[-1<x-3<1\] or \[2<x<4\] which means therefore \[|x+4|<8\]

OpenStudy (anonymous):

so we say at the outset that \[|x-3|<1\] which means \[|(x-3)(x+4)|<8|x-3|\] and we want \[|(x-3)(x+4)|<\epsilon\] \[\delta =\frac{\epsilon}{8}\] and therefore if \[|x-3|<\frac{\epsilon}{8}\]we have making \[|(x-3)(x+4)|<8|(x-3)|<8\times \frac{\epsilon}{8}=\epsilon\]

OpenStudy (anonymous):

wow!! taht was a lot of work!!! is this the end of the problem?

OpenStudy (anonymous):

that is it. all steps are there for you to look at. hope it becomes clear more or less

OpenStudy (anonymous):

YES..THANK U!!! im hopeless when it comes to limits...i do not understand them very well. can u help me with a proof of a limit??

OpenStudy (anonymous):

another one?

OpenStudy (anonymous):

post and i will see what we can do

OpenStudy (anonymous):

okay